Outback Outfitters sells recreational equipment. One of the companys products, a small camp stove, sells for $100 per unit.

4. Refer to the data in (3) above. How many stoves would have to be sold at the new selling price to yield a minimum net operating income of $79,000 per month? (Round your answer to the nearest whole number.)
Answer #1
1.
Contribution margin per unit = Selling price per unit - Variable cost per unit = 100 - 70 30
Contribution margin ratio = Contribution margin per unit / Selling price per unit = 30 / 100 30%
Break-Even Point
Number of stoves = Total Fixed cost / Contribution margin per unit = 144000 / 30 4800
Total sales dollars = Total fixed costs / Contribution margin ratio = 144000 / 30% 480000
2.
If the variable expense per stove increases as a percentage of the selling price and the fixed costs remains unchanged then it will result in higher break-even point, because the contribution margin per unit will decrease which will lead to higher break-even point
3.
Proposed selling price = Present selling price * ( 1 - % reduction ) = 100 * ( 1 - 10% ) 90
Proposed sales of stoves = Present sales of stoves * ( 1 + % increase ) = 17000 * ( 1 + 25% ) 21250
Outback Outfitters
Present Proposed
17000 Stoves 21250 Stoves
Total Per unit Total Per unit
Sales 1700000 100 1912500 90
Variable costs 1190000 70 1487500 70
Contribution margin 510000 30 425000 20
Fixed costs 144000 144000
Net operating income 366000 281000
4.
Number of stoves to be sold = ( Net operating income + Fixed costs ) / Contribution margin per unit = ( 79000 + 144000 ) / 20 11150